if the degree measures of the four angles of a quadrilateral are in the ration 1:2:3:4 what is the degree measure of the largest angle
Answer by stanbon(75887)

if the degree measures of the four angles of a quadrilateral are in the ration 1:2:3:4 what is the degree measure of the largest angle
----
Equation:
x + 2x + 3x + 4x = 360 degrees
10x = 360
x = 36
----
Largest = 4x = 4*36 = 144 degrees
===================================
Cheers,
Stan H.
===============